OpenStudy (anonymous):

Can someone help me with this problem? The atomic mass of element X is 33.97 amu. A 13.33-g sample of X combines with 84.66 g of another element Y to form a compound XY3. Calculate the atomic mass of Y

OpenStudy (anonymous):

first you write hypothetic equation X + 3Y -> XY3 now knowing that you can see that n(XY3) = n(X) so n(X) = m(X) / Ar(X) = 13,33 g / 33,97 gmol-1 and now you can calculate that n=m/M -> M = m/n so M(XY3) = (13,33+84,66)g/n(XY3) and then you can just substract M(XY3) - M(X) = M(Y3) and M(Y) = M(Y3)/3

OpenStudy (anonymous):

I'm sorry I have no idea what you just explained to me :c

OpenStudy (anonymous):

lets say you are going to meet your three friends which are triplets (identical twins wich weih the same), you meet them and you all decide to go and play with a scale, you step on scale and it turns out you weih 10 units of mass, then all of you step on a scale and you get that your mass together is lets say 60. now calculate how much each of the twins weih. thats the line of reasoning behind this task you have. try to calculate this what i gave you and then read your task and see if it becomes more clear.

OpenStudy (anonymous):

I got it!! Thank you very much! :)